OpenStudy (anonymous):

Find the specified nth term in the expansion of the binomial. (Write the expansion in descending powers of x.)

OpenStudy (anonymous):

(x+5y)^10, n=4

OpenStudy (anonymous):

Are you familiar with this ? http://en.wikipedia.org/wiki/Binomial_coefficient

OpenStudy (anonymous):

yes, partially, i'm just learning about it now. i haven't tried that pascal triangle yet, but the problem wants me to use binomial expansions i think

OpenStudy (anonymous):

Here is the definition of binomial expansion: http://en.wikipedia.org/wiki/Binomial_expansion Looks like you are supposed to use pascal's triangle - I'm not aware of another way, but satellite might :)

OpenStudy (anonymous):

@fiddlearound i always get the nth term confused

OpenStudy (anonymous):

the other way is to write out what the coefficients are

OpenStudy (anonymous):

the 4th term here will look like \[\text{some number}\times x^7y^3\]

OpenStudy (anonymous):

and the question is, what is the "some number" you need \[\dbinom{10}{3}\times 5^3\] the \[5^3\] because you really have \[(5y)^3\]

OpenStudy (anonymous):

so your real job is to cube 5 and get 125, and also to find 10 choose 3. do you know how to find that?

OpenStudy (anonymous):

it's 15,000 x^7y^3

OpenStudy (anonymous):

ok but do you know how to get it?

OpenStudy (anonymous):

thanks for helping out!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yeah, i think i understand, i have to move on to the next problem in my school work though, i have a ton to get done

OpenStudy (anonymous):

just for the sake of completeness \[\dbinom{10}{3}=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

yep, in short: \[\left(\begin{matrix}10 \\ 3\end{matrix}\right)\times x ^{10-3} (5y) ^{3}=\frac{10!}{3!*(10-3)!}*125*x^7y^3=120*125*x^{7}y^3=15000x^7y^3\]